OpenStudy (anonymous):

An arc of length 5pi cm and the sector it bounds has an arc of 20pi cm.Find the radius of the circle.

OpenStudy (anonymous):

Area of sector = angle in radians * r^2 * 0.5 20 \pi = angle in radians * r^2 * 0.5 40 \pi / r^2 = angle in radians Length of minor arc = angle in radians * radius (r) 5 \pi / radius = angle in radians So.... 40 \pi / radius^2 = 5 \pi / radius 5 \pi radius^2 / radius = 40 \pi 5 \pi radius = 40 \pi 5(pi)r = 40pi (which you got) r= 40pi/5pi pis cancel out, r=8